# Random coverage from within with variable radii, and Johnson-Mehl cover timesthanks: Supported by EPSRC grant EP/T028653/1

### v2

@@ Appendix A Confirming consistency with Chiu’s result
In this appendix we verify that our Proposition 2.8 is consistent with [1, Theorem 4]. The latter result is concerned with a quantity denoted $T_{L}$ in [1] which is the same as our $\tilde{\tau}_{L}$, assuming we take $v=\gamma=1$ in [1]. Chiu takes $A=[0,1]^{d}$ and defines $c=c(L)=d(d+1)\log L-\log\omega_{d}$. Then in our notation, the result in [1] says that for any $u\in\mathbb{R}$, as $L\to\infty$ we have
$\mathbb{P}\left[c^{d/(d+1)}\omega_{d}^{1/(d+1)}\tilde{\tau}_{L}-c-\log\left(c^{1/(d+1)}\left(\frac{c+\log c}{d+1}\right)^{d-1}\right)\leq u\right]\to F(u),$ (A.1)
@@
To verify that we can recover (A.1) from Proposition 2.8, note first that $c\to\infty$ as $L\to\infty$ and $\log c=\log\log L+\log(d(d+1))+o(1)$. Also
@@
Hence for some $z_{L}$ which tends to zero as $L\to\infty$, the left hand side of (A.1) equals
@@
By (2.12) from Proposition 2.8, and the continuity of the limiting cumulative distribution function in that result, this probability tends to
@@
In other words, we have derived (A.1) from Proposition 2.8.
Acknowledgment. We thank Edward Crane for a useful conversation on the link between the the SPBM and the Johnson-Mehl model.

### v3

@@ Appendix A Confirming consistency with Chiu’s result
In this appendix we verify that our Proposition 2.1 is consistent with [3, Theorem 4]. The latter result is concerned with a quantity denoted $T_{L}$ in [3] which is the same as our $\tilde{\tau}_{L}$, assuming we take $v=\gamma=1$ in [3]. Chiu takes $A=[0,1]^{d}$ and defines $c=c(L)=d(d+1)\log L-\log\omega_{d}$. Then in our notation, the result in [3] says that for any $u\in\mathbb{R}$, as $L\to\infty$ we have
$\mathbb{P}\left[c^{d/(d+1)}\omega_{d}^{1/(d+1)}\tilde{\tau}_{L}-c-\log\left(c^{1/(d+1)}\left(\frac{c+\log c}{d+1}\right)^{d-1}\right)\leq u\right]\to F(u),$ (86)
@@
To verify that we can recover (86) from Proposition 2.1, note first that $c\to\infty$ as $L\to\infty$ and $\log c=\log\log L+\log(d(d+1))+o(1)$. Also
@@
Hence for some $z_{L}$ which tends to zero as $L\to\infty$, the left hand side of (86) equals
@@
By (7) from Proposition 2.1, and the continuity of the limiting cumulative distribution function in that result, this probability tends to
@@
In other words, we have derived (86) from Proposition 2.1.
Acknowledgement. We thank Edward Crane for a useful conversation on the link between the SPBM and the Johnson-Mehl model. We also thank two anonymous referees for some helpful comments.
Data availability statement. The code used to generate Figure 1 and Figure 7 and the videos discussed in Section 2, as well as all the code and data used to create Figure 2, are all available at https://github.com/frankiehiggs/johnson-mehl.